# An Exploratory Case Study on Data Breach Journalism

### v1

@@ 4.3. Sentiments
Sentiment analysis is the final step in the empirical exploration. Thus, Fig. 5 displays the sentiment polarity scores for the news articles and their titles. Unlike what was expected, the news articles split almost perfectly into half between negative and positive polarity scores. Moreover, there are not many articles with values around zero, indicating that the news articles tend to exhibit relatively strong negative or positive sentiments. This observation is even more surprising when consider the short length of the articles. Emotionally loaded language provides one explanation, although it is difficult to speculate why positive sentiments have been common in the articles. In contrast, the articles with negative sentiment polarity scores are presumably at least partially due to the relatively widespread appearance of crime-related topics in the news articles.
@@ 5.1. Conclusion
This paper investigated the previously unexplored topic of data breach journalism with an exploratory case study of databreaches.net. The results can be summarized with the following points:
  • The news site has kept a steady publishing pace; on average, two data breach news articles have appeared on each day after the site’s launch in 2008.
@@
  • The news are often sourced with primary material from the victim organizations. Although also social media is present, additional news material includes high-quality sources, including technology news sites, criminal justice systems, and data protection authorities.
@@ 5.2. Limitations
Finally, the limitations of the sentiment analysis algorithm should be briefly acknowledged. The VADER algorithm was developed with social media analytics in mind. Hence, its lexicon may not be the optimal choice in the present context, which might require a custom dictionary tailored for data breach news. Also the other common limitations apply, including the algorithm’s poor handling of adverbs, irony, sarcasm, context, and comparative sentences (Bordoloi23, ; Denecke23, ). Sentiment analysis algorithms are also sensitive to different pre-processing choices. In the present context these choices include the sampling strategy in particular.

### v2

@@ 4.3. Sentiments
Sentiment analysis is the final step in the empirical exploration. Thus, Fig. 6 displays the sentiment polarity scores for the news articles and their titles. Unlike what was expected and contrary to existing research results (Sinanaj14, ), the news articles split almost perfectly into half between negative and positive polarity scores. Moreover, there are not many articles with values around zero, indicating that the news articles tend to exhibit relatively strong negative or positive sentiments. This observation is even more surprising when consider the short length of the articles. Emotionally loaded language provides one explanation, although it is difficult to speculate why positive sentiments have been common in the articles. In contrast, the articles with negative sentiment polarity scores are presumably at least partially due to the relatively widespread appearance of crime-related topics in the news articles. Moreover, as can be seen from Fig. 7, the sentiment polarity scores have remained stable across time. Thus, the negative or positive sentiments cannot be attributed to a sudden change in the news site’s reporting style.
@@ 5.1. Conclusion
This paper investigated the mostly unexplored topic of data breach journalism with an exploratory case study of databreaches.net. The paper’s results can be summarized with the following points:
  • The news site has kept a steady publishing pace, although there have also been some gaps in the publishing over the years. On average, two data breach news articles have appeared on each day after the site’s launch in 2008.
@@
  • The news are often sourced with primary material from the victim organizations. Although also social media is present, additional news material includes high-quality sources, including technology news sites, criminal justice systems, and data protection authorities. That said, practically all of the sources are in English, which likely reflects a geographic bias.
@@ 5.2. Limitations
Finally, limitations of the sentiment analysis algorithm should be briefly acknowledged. The VADER algorithm was developed with social media analytics in mind. Hence, its lexicon may not be the optimal choice in the present context, which might require a custom dictionary tailored specifically for data breach news. This point generalizes toward other, newer sentiment analysis algorithms. Also the other common limitations apply, including sentiment analysis algorithms’ poor handling of adverbs, irony, sarcasm, context, and comparative sentences (Bordoloi23, ; Denecke23, ). Sentiment analysis algorithms are also sensitive to different pre-processing choices. In the present context these choices include the sampling strategy in particular.